OpenStudy (anonymous):

find the x and y-intercepts of f(x)= |x-6|-12 i'm not sure how to find x because |x-6|-12=? like what am i supposed to equal it to to solve it. and there is no y to find a y intercept, and i don't know how to find it

jimthompson5910 (jim_thompson5910):

f(x)= |x-6|-12 is the same as y = |x-6|-12 since y = f(x)

jimthompson5910 (jim_thompson5910):

to find the x-intercept, plug in y = 0 and solve for x to find the y-intercept, plug in x = 0 and solve for y